OpenStudy (anonymous):

Please Help! Quadratic Formula One number is 5 more than aother number. Find these numbers if their product is -3.

OpenStudy (anonymous):

Let one number be X. Other number = X - 5 Product = (x)(x-5) = -3 thus x^2 - 5x + 3 = 0 solve for x: (x-3)(x-2)=0

OpenStudy (mathstudent55):

Let the smaller number be x The other number is 5 more so it is x + 5 Their product is -3, so \( x (x + 5) = -3 \)

OpenStudy (mathstudent55):

\(x (x + 5) = -3 \) \(x^2 + 5x + 3 = 0 \) This is not factorable, so you need to use the quadratic formula as tye problem states.

OpenStudy (mathstudent55):

^the

OpenStudy (mathstudent55):

Do the quadratic formula again because that's not correct.

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{ b^2 - 4ac}} {2a} \)

OpenStudy (mathstudent55):

\( x = \dfrac{-5 \pm \sqrt{ 5^2 - 4(1)(3)}} {2(1)} \)

OpenStudy (ja1):

Just substitute all your terms into the equation, can you list me the value of each term A,B, and C ?

OpenStudy (ja1):

Nevermind mathstudent got it for you :)

OpenStudy (mathstudent55):

Can you simplify that now?

OpenStudy (ja1):

Remember we use the PEMDAS methodshere like in most algebraic equations.

OpenStudy (ja1):

So start by simplifying the parenthesis :)

OpenStudy (mathstudent55):

@najaas You had two mistakes. 1. the first term in the quadratic formula is -b, so you need -5, not 5 2. inside the root, it's 25 - 12, not 25 + 12

OpenStudy (anonymous):

@mathstudent55 has the right solution. @dauspex you sort of messed up.

OpenStudy (anonymous):

so the answer is -9 and 4?

OpenStudy (mathstudent55):

No. sqrt(13) is irrational and can't be simplified.

OpenStudy (mathstudent55):

\( x = \dfrac{-5 \pm \sqrt{ 25 - 12}} {2}\) \( x = \dfrac{-5 \pm \sqrt{ 13}} {2}\) \( x = \dfrac{-5 + \sqrt{ 13}} {2}\) or \( x = \dfrac{-5 - \sqrt{ 13}} {2}\)

OpenStudy (anonymous):

ooh yeah, okay, I forgot about the square root

OpenStudy (mathstudent55):

Remember, we let x be the smaller of the two numbers. For each of the two solutions, you now add 5 to find the second number. The final solution is two pairs of numbers.

OpenStudy (anonymous):

how do you add -5 and the root of 13 ? @mathstudent55

OpenStudy (mathstudent55):

You must use a common denominator: \( \dfrac{-5 - \sqrt{ 13}} {2} + 5\) \(= \dfrac{-5 - \sqrt{ 13}} {2} + \dfrac{5 \times 2}{2}\) \(= \dfrac{-5 - \sqrt{ 13}} {2} + \dfrac{10}{2}\) \(= \dfrac{5 - \sqrt{ 13}} {2} \)

OpenStudy (mathstudent55):

Then you do the same for the other solution.

OpenStudy (anonymous):

okay, thanks @mathstudent55